OpenStudy (hellomynameiskam):

What is the exact area of a circle with a radius of 31 cm? A. 15.5 cm² B. 31 cm² C. 961 cm² D. 240.25 cm²

OpenStudy (bmk614):

\[\pi*r^2 = \pi*31^2 = 3,019.07\] There is not a choice for a circle with this radius.

OpenStudy (hellomynameiskam):

hmm. i would say its D then idk